Ticket Number: HEA-NOC /20020503-04 Ticket Status: CLOSE
Ticket Type: Unscheduled Ticket Source: HEA-NOC
Ticket Scope: Line Site/Line: HEAnet <==> TEAGASC
Ticket Owner: [log in to unmask] Problem Fixer: HEA-NOC
Ticket Opened: Problem Start: 20020503 1558
Ticket Closed: 20020510 08:33:44 UTC Problem End:
All traffic between Teagasc and HEAnet.
20020503 1505 CW Teagasc confirm nothing happening onsite to cause fault.
20020503 1510 CW Reporting fault to Eircom ticket: 339 Requesting alarm check 1400 - 1515hrs
20020503 1635 CW Eircom report no alarms on line
20020503 2137 CW OSPF session only reestablished at 1811
20020507 0909 CW Hits on line over the weekend 20020506 1326 & 20020507 0238 & 0331. Eircom issue New ticket 408
20020507 0950 CW Eircom see none of these faults, engineer will monitor over the day.
20020507 1050 CW awaiting callback fomr Eircom, additional faults on the circuit
20020507 1132 CW rang Eircom again, still unable to get engineer
20020507 1300 CW Eircom did not see the two outages this morning 0939 & 0941, Engineer is going to inspect HDSL stats and clear counters, They may wish to take the circuit for test overnight.
20020507 1612 CW additional transittion at 1602, awaitng callback
20020507 1724 CW Eircom report errors on Exchange <==> Teagasc trunk, will dispatch a field team to inspect. Circuit transitioned at 1710 again
20020508 1451 CW check on 101 trunk, Report as clear, returned to engineering
20020510 0932 Circuit has been stable for over 24hours. Eircom's inspection of HDSL modem reported as no faults found. Closing ticket
Time to Fix: